OpenStudy (anonymous):

Can anyone explain to me how to prepare 500 ml of buffer containing 3%(w/v) PVP by using PVP powder and Distilled water?I only want the instruction on how to do it...what does (w/v) mean and etc.

OpenStudy (koikkara):

This common usage of % to mean wt/v in biology.

OpenStudy (anonymous):

Then do you know how it can be answered?

OpenStudy (koikkara):

@terenzreignz: @opcode @AriPotta:.... well, thinking.....

OpenStudy (anonymous):

ok

OpenStudy (unklerhaukus):

percentage weight of solute per volume of solution if %(w/v)=x, then there are x grams of solute for every 100 mls of solution

OpenStudy (anonymous):

Ok thanks a lot :)

OpenStudy (anonymous):

Thanks a lot :)

OpenStudy (unklerhaukus):

note that the amount of water in the solution is not necessarily 100 ml

OpenStudy (anonymous):

Then the 100 ml is the normal one,right?

OpenStudy (unklerhaukus):

100 mls is the unit of the solution

OpenStudy (anonymous):

Then at first,I should weight 3 grams of PVP and then add water proportional to have 500 ml,am I right?

OpenStudy (unklerhaukus):

well that would make a 0.6%(w/v) solution

OpenStudy (unklerhaukus):

you need more powder

OpenStudy (anonymous):

How do you calculate?Any formula?Sorry for a lot questions.

OpenStudy (unklerhaukus):

\[3\%(w/v)=\frac{3[\text g]}{100[\text{mls}]}\] \[500[\text{mls}]\cdot3\%(w/v)=500[\text{mls}]\cdot\frac{3[\text g]}{100[\text{mls}]}=...[\text g]\]

OpenStudy (anonymous):

Thanks...really helped ^.^